# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v1

@@ Abstract
Conventional survival metrics, such as Harrell’s concordance index and the Brier Score, rely on the independent censoring assumption for valid inference in the presence of right-censored data. However, when instances are censored for reasons related to the event of interest, this assumption no longer holds, as this kind of dependent censoring biases the marginal survival estimates of popular nonparametric estimators. In this paper, we propose three copula-based metrics to evaluate survival models in the presence of dependent censoring, and design a framework to create realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our empirical analyses in synthetic and semi-synthetic datasets show that our metrics can give error estimates that are closer to the true error, mainly in terms of predictive accuracy.
@@ 1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. However, survival prediction differs as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who are still alive at the end of the study, as well as others who left the study and were “lost to follow-up”. The independent censoring assumption states that, within any subgroup of interest, we assume that the patients (i.e., instances) who are censored at time $t$ are representative of all the patients in that subgroup who remained at risk at time $t$ with respect to their survival experience [Emura and Chen, 2018, Ch. 1]. That is, censoring is random within the subgroups. Let $T_{E}$ be the event distribution and $T_{C}$ be the censoring distribution, then we have $T_{E}\perp T_{C}$i.e.$T_{E}$ provides no information about the distribution of $T_{C}$, and vice versa [Kleinbaum and Klein, 2012, Ch. 1]. If there are features (covariates) $X$, this assumption becomes the conditional independent censoring assumption, i.e. $T_{E}\perp T_{C}\mid X$ [Kleinbaum and Klein, 2012, Ch. 1]. (More details wrt censoring are in Appendix C.) Note that many standard survival tools – e.g., the Kaplan-Meier (KM) [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on an assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1].
@@
Figure 1: (a) Residual dependence between event time $T_{E}$ (cancer relapse) and censoring time $T_{C}$ (cancer death) denoted by the diamond ‘Dep’, due to an unobserved confounding covariate – here ‘Tumor grade’. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a bias estimation on the marginal survival distribution using the Kaplan-Meier (KM) [Kaplan and Meier, 1958] estimator. Note that the two curves are coincide between range $[0,t_{c,i}]$. (c) Scatter plot generated from a standard exponential distribution under a Clayton [Clayton, 1978] copula model.
@@
Figure 1(a) shows how the independent censoring assumption fails to hold by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $T_{E}$ (cancer relapse) and $T_{C}$ (cancer death), the variation in the covariate induces changes in both $T_{E}$ and $T_{C}$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $T_{E}$ and cancer death $T_{C}$, and consequently, $T_{E}$ and $T_{C}$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [Emura and Chen, 2018, Ch. 1]. This phenomenon is called residual dependencei.e.dependent censoring not adjusted by the covariates.
@@
To address such dependency, Figure 1(c) presents a copula-based modeling approach. A copula function $C_{\theta}(u,v)$ links random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. Applied here, it captures the dependency between event and censoring distributions ($T_{E}$ and $T_{C}$) by sampling the marginals from a (survival) copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring).
In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring. However, it is typically impossible to verify this in practice because we only observe one outcome (either event or censoring) per instance, but not both [Tsiatis, 1975]. Although we can reduce dependent censoring simply by collecting more data (i.e., features) [Emura and Chen, 2018, Ch. 1], we would still have to evaluate our model under dependent censoring, independent of the learned model, preferably. Popular metrics for evaluating predicted survival probabilities (or risk scores) are Harrell’s concordance index (CI) [Harrell Jr et al., 1982], Uno’s CI [Uno et al., 2011], the Brier Score (BS) [Brier, 1950], and the mean absolute error (MAE) using one of hinge [Haider et al., 2020], margin [Haider et al., 2020] or pseudo times [Qi et al., 2023a]. Unfortunately, dependent censoring is both common, and also introduces a bias that can affect the estimation of survival probabilities, and none of these metrics account for this dependency. Alternative methods have been proposed to handle dependent censoring, such as the inverse probability of censoring weighting (IPCW) [Robins and Finkelstein, 2000], but this approach still uses the KM estimator. Motivated by these limitations, we introduce a new approach to evaluating survival models under dependent censoring. To our knowledge, no prior work has proposed evaluation metrics that explicitly model the degree of dependent censoring between $T_{E}$ and $T_{C}$. Our main contributions are:
  • We show that conventional survival metrics, such as Harrell’s CI, Uno’s CI, the IPCW-integrated Brier Score, and the marginal MAE, are biased under dependent censoring, meaning they will give inaccurate estimates of the model’s discriminative or predictive performance.
  • We propose three novel metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These metrics can evaluate the concordance (i.e., ranking), the accuracy of the survival curve, and the accuracy of the predicted time-to-event estimates, respectively, by considering the dependency between the censoring and event distributions.
@@
  • We provide empirical evidence that our proposed metrics can provide more accurate error measures for survival models, mainly in terms of predictive accuracy, i.e., Brier Score and MAE.
@@ 2.1 Survival data and notation
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [Kuan et al., 2023], predicting the time to a fall in older adults [Lillelund et al., 2024a], or evaluating a new experimental treatment for cancer [Zhang et al., 2011]. Most often, the observation for a given instance will consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S(t)=1-F(t)=\text{Pr}(T>t)$. Note that $S(t)$ (resp., $F(t)$) correspond to the probabilities of being event-free (resp., having experienced the event by) time $t$. We can estimate $S\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [Kaplan and Meier, 1958] estimator (see Appendix H for a detailed derivation).
@@
If features are entered into the model, popular methods include the CoxPH model [Cox, 1972, Katzman et al., 2018, Lillelund et al., 2024b], gradient-boosted regression [Ridgeway, 1999], Random Survival Forest [Ishwaran et al., 2008] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. In this context, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is the “censoring bit”: If $\delta_{i}=0$, then $t_{i}=t_{c,i}$, where $t_{c,i}$ is the time of censoring, otherwise, if $\delta_{i}=1$, then $t_{i}=t_{e,i}$, where $t_{e,i}$ is the time of the event. Given survival data $\mathcal{D}$, let $f_{T\mid X}(\cdot)$ and $F_{T\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon (i.e., from 00 to $t_{\text{max}}$). Given the covariates, we then have the following useful definitions:
@@
$h_{T\mid X}(t\mid X)$ =\ = $\lim_{\epsilon\to 0}\text{Pr}(T\in[t,t+\epsilon)\mid T\geq t,X)$ (2.3)
=\ = $\frac{f_{T\mid X}(t\mid X)}{S_{T\mid X}(t\mid X)}$
@@ 2.2 Copula-Based Estimation
Zheng and Klein [1995] generalized the KM estimator under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [Emura and Chen, 2016, Emura and Michimae, 2017, Hoora Moradian and Bellavance, 2019]. Rivest and Wells [2001] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). Given survival data without covariates, i.e.$(t_{i},\delta_{i})$ for $i=1,\ldots,n$, where $t_{i}=\min(t_{e,i},t_{c,i}),\;\delta_{i}=I(t_{e,i}\leq t_{c,i})$ with $I(\cdot)$ denotes the indicator function, we can estimate the survival function using the CG estimator (see Appendix H for a detailed derivation):
@@
$\small\begin{split}\hat{S}_{T}(t)&=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\\ \hbox{for}&\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.4)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{T}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{T}(t)$ is undefined for $t>t_{\text{max}}$ [Emura and Chen, 2018, Ch. 4].
@@ 2.3 Evaluation metrics and limitations
          $\text{Pr}(\delta_{i}=1,t_{i}<t_{j})\quad\neq\quad\text{Pr}(t_{e,i}<t_{e,j}).$
@@
$\begin{array}[]{l}\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})\\ =\quad\text{Pr}(\,t_{e,i}<t_{c,i},\ \ t_{e,i}\ <\ \min\{t_{c,j},t_{e,j}\}\ )\\ =\quad\text{Pr}(\,t_{e,i}<t_{c,i},\ t_{e,i}<t_{c,j},\ t_{e,i}<t_{e,j})\\ =\quad\underline{\text{Pr}(\,t_{e,i}<t_{c,i})\ \text{Pr}(t_{e,i}<t_{c,j})}\quad\text{Pr}(t_{e,i}<t_{e,j})\ ,\end{array}$
@@
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [2020] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $t_{i}$, its margin time is calculated as:
@@
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption and its computational simplicity – requiring no interaction with covariates – it becomes problematic when censoring and event times are dependent [Emura and Chen, 2018, Ch. 1]. In such cases, the KM estimates may be biased and inconsistent [Campigotto and Weller, 2014]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [Gerds and Schumacher, 2006]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.

### v2

@@ Abstract
Conventional survival metrics, such as Harrell’s concordance index and the Brier Score, rely on the independent censoring assumption for valid inference in the presence of right-censored data. However, when instances are censored for reasons related to the event of interest, this assumption no longer holds, as this kind of dependent censoring biases the marginal survival estimates of popular nonparametric estimators. In this paper, we propose three copula-based metrics to evaluate survival models in the presence of dependent censoring, and design a framework to create realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our empirical analyses in synthetic and semi-synthetic datasets show that our metrics can give error estimates that are closer to the true error, mainly in terms of prediction accuracy.
@@ 1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. However, survival prediction differs as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who are still alive at the end of the study, as well as others who left the study and were “lost to follow-up”. The independent censoring assumption states that, within any subgroup of interest, we assume that the patients (i.e., instances) who are censored at time $t$ are representative of all the patients in that subgroup who remained at risk at time $t$ with respect to their survival experience [Emura and Chen, 2018, Ch. 1]. That is, censoring is random within the subgroups. Let $E$ be the event distribution and $C$ be the censoring distribution, then we have $E\perp C$i.e.$E$ provides no information about the distribution of ${C}$, and vice versa [Kleinbaum and Klein, 2012, Ch. 1]. If there are features (covariates) $\bm{X}$, this assumption becomes the conditional independent censoring assumption, i.e. $E\perp C\mid\bm{X}$ [Kleinbaum and Klein, 2012, Ch. 1]. (See Appendix C for details.) Note that many standard survival tools – e.g., the Kaplan-Meier (KM) [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on an assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1].
@@
Figure 1: (a) Residual dependence between event time $E$ (cancer relapse) and censoring time $C$ (cancer death) denoted by the diamond ‘Dep’, due to an unobserved confounding covariate – here ‘Tumor grade’. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a bias estimation on the marginal survival distribution using the Kaplan-Meier (KM) [Kaplan and Meier, 1958] estimator. Note that the two curves are coincide between range $[0,c_{i}]$. (c) Scatter plot generated from a standard exponential distribution under a Clayton [Clayton, 1978] copula model.
@@
Figure 1(a) shows how the independent censoring assumption fails to hold by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $E$ (cancer relapse) and $C$ (cancer death), the variation in the covariate induces changes in both $E$ and $C$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $E$ and cancer death $C$, and consequently, $E$ and $C$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [Emura and Chen, 2018, Ch. 1]. This phenomenon is called residual dependencei.e.dependent censoring not adjusted by the covariates.
@@
To address such dependency, Figure 1(c) presents a copula-based modeling approach. A copula function $C_{\theta}(u,v)$ links random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1] and have recently gained interest in survival analysis [Foomani et al., 2023, Zhang et al., 2024, Liu et al., 2024]. Applied here, it captures the dependency between event and censoring distributions by sampling the marginals from a (survival) copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring).
In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring. However, it is typically impossible to verify this in practice because we only observe one outcome (either event or censoring) per instance, but not both [Tsiatis, 1975]. Although we can reduce dependent censoring simply by collecting more data (i.e., features) [Emura and Chen, 2018, Ch. 1], we would still have to evaluate our model under dependent censoring, independent of the learned model, preferably. Popular metrics for evaluating predicted survival probabilities (or risk scores) are Harrell’s concordance index (CI) [Harrell Jr et al., 1982], the Brier Score (BS) [Brier, 1950], and the mean absolute error (MAE) using one of hinge [Haider et al., 2020], margin [Haider et al., 2020] or pseudo times [Qi et al., 2023a]. Unfortunately, dependent censoring is both common [Emura and Chen, 2018, Ch. 1] and also introduces a bias that can affect the estimation of survival probabilities, and none of these metrics account for this. Alternative methods have been proposed to handle dependent censoring, such as Uno’s CI [Uno et al., 2011] using the inverse probability of censoring weighting (IPCW) [Robins and Finkelstein, 2000], but this approach still relies on the KM estimator. Motivated by these limitations, we introduce three new metrics for evaluating survival models under dependent censoring. To our knowledge, no prior work has proposed evaluation metrics that explicitly model the degree of dependency between $E$ and $C$. Our main contributions are:
  • We show that conventional survival metrics, such as Harrell’s CI, Uno’s CI, the IPCW-integrated Brier Score, and the marginal MAE, are biased under dependent censoring, meaning they can give inaccurate estimates of a model’s performance.
  • We propose three novel metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These metrics can evaluate the concordance (i.e., ranking), the accuracy of the survival curve, and the accuracy of the predicted time-to-event estimates, respectively, by considering the dependency between event and censoring.
@@
  • We provide empirical evidence that our proposed metrics can provide more accurate error measures for survival models, mainly in terms of prediction accuracy.
@@ 2.1 Survival data and notation
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [Kuan et al., 2023], predicting the time to a fall in older adults [Lillelund et al., 2024a], or evaluating a new experimental treatment for cancer [Zhang et al., 2011]. Most often, the observation for a given instance will consist of a random variable, $E$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $E$ can be characterized by the probability distribution function $F_{E}(t)=\text{Pr}(E\leq t)$ or, equivalently, by the survival function $S_{E}(t)=1-F_{E}(t)=\text{Pr}(E>t)$. Note that $S_{E}(t)$ (resp., $F_{E}(t)$) correspond to the probabilities of being event-free (resp., having experienced the event by) time $t$. We can estimate $S_{E}\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [Kaplan and Meier, 1958] estimator (see Appendix H for a detailed derivation).
@@
If features are entered into the model, popular methods include the CoxPH model [Cox, 1972, Katzman et al., 2018, Lillelund et al., 2024b], gradient-boosted regression [Ridgeway, 1999], Random Survival Forest [Ishwaran et al., 2008] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. In this context, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is a event indicator where $\delta_{i}=0$ means the event has not happened by $t_{i}$ (right-censored) and $\delta_{i}=1$ means the event occurred at $t_{i}$ (uncensored). We consider each instance to have a potential event time $e_{i}$ and a right-censoring time $c_{i}$, where we observe only the earlier of the two. Therefore, we define the observed time and event indicator as:
          $t_{i}\ :=\ \min\{e_{i},c_{i}\},\quad\text{and}\quad\delta_{i}\ :=\ \mathbbm{1}[e_{i}\leq c_{i}],$
where $\mathbbm{1}[\cdot]$ represents the indicator function.
Given survival data $\mathcal{D}$, let $f_{E\mid\bm{X}}(\cdot)$ and $F_{E\mid\bm{X}}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon (i.e., from 00 to $t_{\text{max}}$). With these two functions, we then have the following useful definitions:
@@
$h_{E\mid\bm{X}}(t\mid\bm{X})$ =\ = $\lim_{\epsilon\to 0}\text{Pr}(E\in[t,t+\epsilon)\mid E\geq t,\bm{X})$ (2.3)
=\ = $\frac{f_{E\mid\bm{X}}(t\mid\bm{X})}{S_{E\mid\bm{X}}(t\mid\bm{X})}$
@@ 2.2 Copula-Based Estimation
Zheng and Klein [1995] generalized the KM estimator under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [Emura and Chen, 2016, Emura and Michimae, 2017, Hoora Moradian and Bellavance, 2019]. Rivest and Wells [2001] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). We can estimate the survival function using the CG estimator (see Appendix H for a detailed derivation):
@@
$\small\begin{split}\hat{S}_{E}(t)&=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\\ \hbox{for}&\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.4)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{E}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{E}(t)$ is undefined for $t>t_{\text{max}}$ [Emura and Chen, 2018, Ch. 4].
@@ 2.3 Evaluation metrics and limitations
          $\text{Pr}(\delta_{i}=1,t_{i}<t_{j})\quad\neq\quad\text{Pr}(e_{i}<e_{j}).$
@@
$\begin{array}[]{l}\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})\\ =\quad\text{Pr}(\,e_{i}<c_{i},\ \ e_{i}\ <\ \min\{c_{j},e_{j}\}\ )\\ =\quad\text{Pr}(\,e_{i}<c_{i},\ t_{e,i}<c_{j},\ e_{i}<e_{j})\\ =\quad\underline{\text{Pr}(\,c_{i}<c_{i})\ \text{Pr}(e_{i}<c_{j})}\quad\text{Pr}(e_{i}<e_{j})\ ,\end{array}$
@@
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [2020] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $c_{i}$, its margin time is calculated as:
@@
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption and its computational simplicity – requiring no interaction with the covariates – it becomes problematic when censoring and event times are dependent [Emura and Chen, 2018, Ch. 1]. In such cases, the KM estimates may be biased and inconsistent [Campigotto and Weller, 2014]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [Gerds and Schumacher, 2006]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.